OpenStudy (anonymous):

how can u rewrite these logarithmic equations an equivalent exponent equation or the other way around m=loga x is equivalent to am=x1. log3 5=y,2. loga 7=-2then solvelog2x=-3,logx9=1/2

OpenStudy (anonymous):

need help solving and rewriting

OpenStudy (anonymous):

can u go over with this with me and help rewrite and solve

OpenStudy (anonymous):

with attachment to help

OpenStudy (anonymous):

Sorry for the delay, my paid gig had some one show up. You've seen: \[\log_{a} b=x \rightarrow a^{x}=b\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\log_{3} 5=y\] What is the "base"?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

No. In exponential form \[3^y=5\] so the 3 is the base

OpenStudy (anonymous):

\[base^{exponent}=answer\]

OpenStudy (anonymous):

In log form \[\log_{base}answer =exponent \]

OpenStudy (anonymous):

OK?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is \[\log_{4} 16 = 2\] in exponential form?

OpenStudy (anonymous):

4y=16,

OpenStudy (anonymous):

right, 4^2 = 16

OpenStudy (anonymous):

If we had to write \[3^2=9\] in log form, we'd write \[\log_{3}9=2 \]

OpenStudy (anonymous):

OK?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Your second problem is \[\log_{a}7=2 \] What is the base here?

OpenStudy (anonymous):

7y=-2

OpenStudy (anonymous):

The base is a (it is that small sub scripted (a)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The exponent is on the right side of the equal sign

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

a7

OpenStudy (anonymous):

Right, so try writing the a, the 7 and the -2 in exponential form

OpenStudy (anonymous):

7a=-2

OpenStudy (anonymous):

a is the base -2 is the exponent and the 7 is the "answer"

OpenStudy (anonymous):

\[a^{-2}=7\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would bea-2=7

OpenStudy (anonymous):

How about \[\log_{a} 30 = 5\] What is that in exponential form?

OpenStudy (anonymous):

Base is ___ Exponent is ____ and the answer is _____

OpenStudy (anonymous):

a30=5

OpenStudy (anonymous):

30

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

right, 30 is the answer a is the base and 5 is the exponent

OpenStudy (anonymous):

Try watching the first few minutes of this video http://patrickjmt.com/properties-of-logarithms-part-1/

OpenStudy (anonymous):

How did the video help?

OpenStudy (anonymous):

ok